OpenStudy (anonymous):

A bullet is shot from the moon at an angle of 41 degrees relative to the horizon, with an initial velocity of 205 m/s. Ignoring curvature of the moon, calculate the total distance the bullet will travel before returning to height=0. (Assume the initial launch height is 0). Please help..never been good at this kind of stuff.

OpenStudy (anonymous):

Moon has gravity 1/6 of Earth's so g' = g/6 = (9.8/6) m/s^2 Will reach peak height when velocity =0 0 = 205 sin(41) - (g')t t = 205 sin(41)/(9.8/6) coming down takes the same t for total of 2t horizontal distance will be (2t) 205 cos (41)

OpenStudy (anonymous):

I hate the cos trig stuff! Thanks

OpenStudy (anonymous):

Practice may not make you love it, but you will get more comfortable with it in time.